# Real-Time Sampling-Based Safe Motion Planning for Robotic Manipulators in Dynamic Environments

### v1

@@ VI Experiments
VI-A Experimental Setup and Implementation Details
@@
[…]
VI-B Obtained Results
@@
To validate the proposed motion planning method in dynamic environments, we conduct four real scenarios (video available here). Representative snapshots are shown in Figs. 20, 21, 22 and 23, where the path traversed by the end-effector is depicted by red lines. The exception is the cyan line that denotes the part of the trajectory before the goal is reached. Time instances $t$ in $\mathrm{[s]}$ are marked in the top part of each figure. The first three scenarios test the regular variant of DRGBT, while the last one deals with DRGBT-safe (from Sec. IV).
@@
Scenario 1 – “Moving-car with balloon” is shown in Fig. 20, and considers a car with the balloon circulating counterclockwise around the robot with a constant speed of cca. $0.36\,\mathrm{[\frac{m}{s}]}$ (i.e., the circle with a radius of $58\,\mathrm{[cm]}$ is completed within $10\,\mathrm{[s]}$). For the start and goal we choose $\boldsymbol{q}_{start}=\left[-\pi,0,-\frac{\pi}{2},0,0,0\right]^{T}$ and $\boldsymbol{q}_{goal}=\left[\pi,0,0,\pi,\frac{\pi}{2},0\right]^{T}$, respectively, and they are swapped when the goal is reached thus emulating repetitive pick-and-place operations. The robot follows an initially planned path during the time $t\in[0,2]$, and then it decelerates in order not to hit the obstacle, after which the replanning is triggered. Similar situation occurs for $t\in[4,5]$ and $t\in[5,7]$. After $10\,\mathrm{[s]}$, the goal is reached. During $t\in[10,13.5]$, the robot and obstacle approach each other. However, the robot successfully avoids the obstacle within $t\in[13.5,16]$. Afterward, the path is replanned, and the goal is therefore reached since the predefined path remains collision-free. The measured joint velocities (for the part of the experiment) are given in Fig. 19. Clearly, they remain within the set limits of $1.5\,\mathrm{[\frac{rad}{s}]}$ for each joint. Analogous diagrams are obtained for other scenarios, yet those are omitted for the limited space.
Figure 19: Measured velocities for each joint from Scenario 1
Figure 20: Snapshots from Scenario 1 – “Moving-car with balloon”
Figure 21: Snapshots from Scenario 2 – “Moving-boxes”
Figure 22: Snapshots from Scenario 3 – “Human as an obstacle”
@@
Figure 23: Snapshots from Scenario 4 – “Safe collaboration/coexistence with human”
@@
Scenario 3 – “Human as an obstacle” from Fig. 22 and Scenario 4 – “Safe collaboration/coexistence with human” from Fig. 23 represent a mock-up of scenario where the robot performs pick-and-place operations while the human operator interrupts the robot by frequent intrusions into the workspace. The robot has to pick up objects from the table while avoiding the human at the same time. To facilitate target object recognition, the robot is programmed to pick only relatively small objects while those larger ones are considered dynamic obstacles. The maximally expected obstacle velocity in Scenario 4 is set to $v_{obs}=0.5\,\mathrm{[\frac{m}{s}]}$, which is particularly relevant to generating dynamic expanded bubbles, while conducted scenarios with $v_{obs}=1.6\,\mathrm{[\frac{m}{s}]}$ can be seen in the accompanying video.
Generally, when obstacles move more swiftly as in Scenarios 3 and 4, the replanning process is triggered more frequently. As for DRGBT regular variant, the robot always avoids the human (e.g., for $t\in[8,10.5]$, $t\in[23,25]$, and $t\in[47.5,49]$ in Fig. 22), and then tries to reach the goal configuration when no risk of collision is observed within a certain range (e.g., for $t\in[10.5,18]$, $t\in[25,27]$, $t\in[27,42]$, and $t\in[49,52]$ in Fig. 22). On the other hand, DRGBT-safe prioritizes the deceleration of the robot in order not to collide with the human. In case a collision occurs, it will be at zero speed of the robot (type 2 collision). Interestingly, if the distance-obstacles proximity remains relatively small (or zero), the robot will not be able to move fast (or to move at all) ensuring human safety (e.g., for $t\in[12.5,14]$, $t\in[26,30]$, and $t\in[60,66]$ in Fig. 23). The robot remains considerably slow until the distance to obstacles increases. It is worth mentioning the robot could not stop timely in $t=60$ since the estimated human arm velocity reached nearly $1.8\,\mathrm{[\frac{m}{s}]}$, which is substantially higher than the assumed limit of $v_{obs}=0.5\,\mathrm{[\frac{m}{s}]}$. Nevertheless, video reveals that increasing $v_{obs}$ up to $1.6\,\mathrm{[\frac{m}{s}]}$ does not yield collisions. Indeed, both variants of DRGBT successfully complete the required pick-and-place task clearly revealing the trade-off between performance and safety.
VII Discussion and Conclusions
@@
Future work directions include a generalization of the used approach to other types of robots (particularly those with more than six DoFs), as well as explicit handling of non-convex obstacles. It would be interesting to investigate the influence of enhanced perception through the use of dedicated hardware (e.g., a separate PC for processing). Finally, tighter integration with the speed and separation monitoring paradigm is desirable for human-robot coexistence contexts.

### v2

@@ VII Experiments
VII-A Experimental Setup and Implementation Details
@@
The planning algorithm is implemented within the “Planning” block [^5] The algorithm is run on the laptop PC with Intel®  Core™  i7-9750H CPU @ 2.60 GHz $\times$ 12 with 16 GB of RAM, with the code compiled to run on a single core of the CPU without any GPU acceleration. The affordability of systems such as Intel NUC PCs, commonly found in robotics applications, give us the idea of what can be considered as a low-cost x86 CPU. In experiments, we use the same PC within the “Processing” block, i.e., $\mathrm{PC1\equiv PC2}$. Generally, they can be different. using ROS2 environment with the open-source code available online here. Besides standard inputs required by the planning, its output is a desired trajectory containing values for $\boldsymbol{q}_{d}$, $\dot{\boldsymbol{q}}_{d}$ and $\ddot{\boldsymbol{q}}_{d}$ of each robot’s joint. Although the planning generates trajectories with a frequency $f_{alg}$, which is usually lower than $f_{con}$, the obtained trajectory can be sampled with $f_{con}$, and then fed to the controller with the corresponding samples in order to achieve a smooth motion of the robot. Since $f_{cam}$ may be lower than $f_{alg}$, it is questionable whether setting $f_{alg}>f_{cam}$ makes sense in practical applications. Theoretically, DRGBT can execute in such cases, since the last available camera measurements can always be used. However, it turns out that using $f_{alg}\gg f_{cam}$ (e.g., $f_{alg}\geq 3f_{cam}$) is not recommended since the algorithm may falsely perceive the obstacles as stationary for a few iterations, thus causing generated plans to be inadequate. This might be addressed by predicting the motion of obstacles (e.g., [67, 68]), however, this exceeds the scope of this paper. Since the simulation study suggests the sweet spot to be around $f_{alg}^{*}=20\,\mathrm{[Hz]}$, the cameras’ frame rate is also set to $f_{cam}=20\,\mathrm{[Hz]}$ for all conducted scenarios in the sequel. Another reason for setting such a value of $f_{cam}$ is achieving a hard real-time execution of the “Processing” block w.r.t. specified voxel grid resolution.
VII-B Obtained Results
@@
To validate the proposed motion planning method in dynamic environments, we conduct four real scenarios (video available here). Representative snapshots are shown in Figs. 21, 22, 23 and 24, where the path traversed by the end-effector is depicted by red lines. The exception is the cyan line that denotes the part of the trajectory before the goal is reached. Time instances $t$ in $\mathrm{[s]}$ are marked in the top part of each figure. The first three scenarios test the regular variant of DRGBT, while the last one deals with DRGBT-safe (from Sec. IV).
@@
Scenario 1 – “Moving-car with balloon” is shown in Fig. 21, and considers a car with the balloon circulating counterclockwise around the robot with a constant speed of cca. $0.36\,\mathrm{[\frac{m}{s}]}$ (i.e., the circle with a radius of $58\,\mathrm{[cm]}$ is completed within $10\,\mathrm{[s]}$). For the start and goal we choose $\boldsymbol{q}_{start}=\left[-\pi,0,-\frac{\pi}{2},0,0,0\right]^{T}$ and $\boldsymbol{q}_{goal}=\left[\pi,0,0,\pi,\frac{\pi}{2},0\right]^{T}$, respectively, and they are swapped when the goal is reached thus emulating repetitive pick-and-place operations. The robot follows an initially planned path during the time $t\in[0,2]$, and then it decelerates in order not to hit the obstacle, after which the replanning is triggered. Similar situation occurs for $t\in[4,5]$ and $t\in[5,7]$. After $10\,\mathrm{[s]}$, the goal is reached. During $t\in[10,13.5]$, the robot and obstacle approach each other. However, the robot successfully avoids the obstacle within $t\in[13.5,16]$. Afterward, the path is replanned, and the goal is therefore reached since the predefined path remains collision-free. The measured joint velocities for Scenario 1 are given in Fig. 20. Clearly, they remain within the set limits of $1.5\,\mathrm{[\frac{rad}{s}]}$ for each joint. Analogous diagrams are obtained for other scenarios, yet those are omitted for the limited space.
Figure 20: Measured velocities for each joint from Scenario 1.
Figure 21: Snapshots from Scenario 1 – “Moving-car with balloon”.
Figure 22: Snapshots from Scenario 2 – “Moving-boxes”.
Figure 23: Snapshots from Scenario 3 – “Human as an obstacle”.
@@
Figure 24: Snapshots from Scenario 4 – “Safe collaboration/coexistence with human”.
@@
Scenario 3 – “Human as an obstacle” from Fig. 23 and Scenario 4 – “Safe collaboration/coexistence with human” from Fig. 24 represent a mock-up of scenario where the robot performs pick-and-place operations while the human operator interrupts the robot by frequent intrusions into the workspace. The robot has to pick up objects from the table while avoiding the human at the same time. To facilitate target object recognition, the robot is programmed to pick only relatively small objects while those larger ones are considered dynamic obstacles. The maximally expected obstacle velocity in Scenario 4 is set to $v_{obs}=0.5\,\mathrm{[\frac{m}{s}]}$, which is particularly relevant to generating $\mathcal{DEB}$s, while conducted scenarios with $v_{obs}=1.6\,\mathrm{[\frac{m}{s}]}$ can be seen in the accompanying video.
Generally, when obstacles move more swiftly as in Scenarios 3 and 4, the replanning process is triggered more frequently. As for DRGBT regular variant, the robot always avoids the human (e.g., for $t\in[8,10.5]$, $t\in[23,25]$, and $t\in[47.5,49]$ in Fig. 23), and then tries to reach the goal configuration when no risk of collision is observed within a certain range (e.g., for $t\in[10.5,18]$, $t\in[25,27]$, $t\in[27,42]$, and $t\in[49,52]$ in Fig. 23). On the other hand, DRGBT-safe prioritizes the deceleration of the robot in order not to collide with the human. In case a collision occurs, it will be at zero speed of the robot (type II collision). Interestingly, if the distance-obstacles proximity remains relatively small (or zero), the robot will not be able to move fast (or to move at all) ensuring human safety (e.g., for $t\in[12.5,14]$, $t\in[26,30]$, and $t\in[60,66]$ in Fig. 24). The robot remains considerably slow until the distance to obstacles increases. It is worth mentioning the robot could not stop timely in $t=60$ since the estimated human arm velocity reached nearly $1.8\,\mathrm{[\frac{m}{s}]}$, which is substantially higher than the assumed limit of $v_{obs}=0.5\,\mathrm{[\frac{m}{s}]}$. Nevertheless, video reveals that increasing $v_{obs}$ up to $1.6\,\mathrm{[\frac{m}{s}]}$ does not yield collisions. Indeed, both variants of DRGBT successfully complete the required pick-and-place task clearly revealing the trade-off between performance and safety.
VIII Discussion and Conclusions
@@
Future work directions include explicit handling of non-convex obstacles. It would be interesting to investigate the influence of enhanced perception through the use of dedicated hardware (e.g., a separate PC for processing). Finally, tighter integration with the speed and separation monitoring paradigm is desirable for human-robot coexistence contexts.